# SANGRIA: Surgical Video Scene Graph Optimization for Surgical Workflow Prediction

### v1

@@ 2.1 Dynamic Scene Graph Generation
For an input image , DINO key features are obtained by partitioning into patches and passing them to DINO. The adjacency matrix is then generated by patchwise dot product as follows:
@@
DSG Optimization We establish a probabilistic estimation of edge weights within the DSGs via equation , where refers to the edge weights between the clusters of DSG, in which indicating the strength of the relations between clusters , . represents the set of trainable parameters. The probabilistic setup allows for flexibility in optimizing the edge weights while learning the downstream tasks and equips the DSG generation to account for the inherent uncertainty and variability present in the unsupervised graph clustering results, leading to more robust and accurate inference.
@@ 2.2 End-to-end Pipeline
Prototype Matching To assign semantic classes to DSG nodes, we create prototypes by leveraging DINO patch features and ground truth (GT) segmentation annotations. Exploiting only 5 annotations per class, patch features corresponding to GT segmentation masks are used as prototype patches. We use the mean of the patch features as a prototype for that object class. To predict the semantic category of a node (cluster), a pairwise cosine similarity with prototypes is calculated. Since different clusters might represent the same region, such as surgical tape, we used argmax to assign class labels to clusters.
@@ 3 Implementation Details
Datasets We experiment on 3 datasets: CATARACTS [3, 2] consists of 50 cataract surgery videos of pixels at fps. The dataset is split 25-5-20 for training, validation, and testing, with videos annotated on 19 surgical phases.
@@ 4 Results & Discussion
Surgical Workflow Prediction
section 4 presents an ablation study on window size and spatial and temporal embeddings as well as a comparative analysis of our proposed method against existing techniques on phase segmentation tasks for CATARACTS and Cataract101 datasets. We show that increasing window size together with temporal embeddings improves phase segmentation performance, while spatial embeddings have minimal impact.
Our method demonstrates superior performance in terms of accuracy and F1 score by indicating accuracy and F1 score improvement over previous graph-based phase segmentation SOTA [11].
As there are no other benchmarks on the CATARACTS dataset, we add an additional non-graph-based solution representing the best practices in the surgical workflow. It exploits DINO as a strong spatial feature extractor followed by long-range temporal learning via TCN++[14]. Our dynamic end-to-end setup outperforms the CNN baseline with accuracy and F1 score.
Our method consistently yields closely competitive results on Cataract101. This demonstrates the robustness and adaptability of our approach, providing a solid foundation for further refinement and application in diverse contexts.
@@
DSG Generation
Figure 3 B manifests improvement of joint optimization of DSG generation with phase segmentation on surgical tool segmentation. The jointly optimized model classifies the "Primary Knife" better than the per frame optimized model. We hypothesize that guidance towards the downstream task significantly improves attention to tools with high importance in the current surgical phase. Our DSG also indicates better explainability by highlighting the importance of graph edges in the end task.
@@ 5 Conclusion
5.0.1 Acknowledgements
[…]
Appendix 0.A SANGRIA: Supplementary Material

### v2

@@ 2.1 Dynamic Scene Graph Generation
For an input image , DINO[5] key features are obtained by partitioning into patches and passing them to DINO. The adjacency matrix is then generated by patchwise dot product as follows:
@@
DSG Optimization We establish a relaxed estimation of edge weights within the DSGs via equation , where refers to the edge weights between the clusters of DSG, in which indicating the strength of the relations between clusters , . represents the set of trainable parameters. The relaxed setup allows for flexibility in optimizing the edge weights while learning the downstream tasks and equips the DSG generation to account for the inherent uncertainty and variability present in the unsupervised graph clustering results, leading to more robust and accurate inference.
@@ 2.2 End-to-end Pipeline
Prototype Matching To assign semantic classes to DSG nodes, we create prototypes [15] by leveraging DINO patch features and ground truth (GT) segmentation annotations. Exploiting only 5 annotations per class, patch features corresponding to GT segmentation masks are used as prototype patches. We use the mean of the patch features as a prototype for that object class. To predict the semantic category of a node (cluster), a pairwise cosine similarity with prototypes is calculated. Since different clusters might represent the same region, such as surgical tape, we used argmax to assign class labels to clusters.
@@ 3 Implementation Details
Datasets We experiment on 3 datasets: CATARACTS [2] consists of 50 cataract surgery videos of pixels at fps. The dataset is split 25-5-20 for training, validation, and testing, with videos annotated on 19 surgical phases.
@@ 4 Results & Discussion
Surgical Workflow Prediction section 4 presents an ablation study on window size and spatial and temporal embeddings as well as a comparative analysis of our proposed method against existing techniques on phase segmentation tasks for CATARACTS and Cataract101 datasets. We show that increasing window size together with temporal embeddings improves phase segmentation performance, while spatial embeddings have minimal impact. Our method demonstrates superior performance in terms of accuracy and F1 score by indicating accuracy and F1 score improvement over previous graph-based phase segmentation SOTA [10]. As there are no other benchmarks on the CATARACTS dataset, we add an additional non-graph-based solution representing the best practices in the surgical workflow. It exploits DINO as a strong spatial feature extractor followed by long-range temporal learning via TCN++[13]. Our dynamic end-to-end setup outperforms the CNN baseline with accuracy and F1 score. Our method consistently yields closely competitive results on Cataract101. This demonstrates the robustness and adaptability of our approach, providing a solid foundation for further refinement and application in diverse contexts.
@@
DSG Generation Figure 3 B manifests improvement of joint optimization of DSG generation with phase segmentation on surgical tool segmentation. The jointly optimized model classifies the "Primary Knife" better than the per frame optimized model. We hypothesize that guidance towards the downstream task significantly improves attention to tools with high importance in the current surgical phase. Our DSG also indicates better explainability by highlighting the importance of graph edges in the end task.
@@ 5 Conclusion
6 SANGRIA: Supplementary Material